# TSINR: Capturing Temporal Continuity via Implicit Neural Representations for Time Series Anomaly Detection

### v2

@@ Abstract.
conference: the 31th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 3–7, 2025; Toronto, ON, Canadaccs: Computing methodologies Neural networksccs: Mathematics of computing Time series analysis
@@ 1. Introduction
Time series anomaly detection, which aims to identify unusual patterns or events across a sequence of data points collected over time (Li and Jung, 2023), has attracted a lot of attention recently. In many fields (e.g., finance, healthcare, manufacturing, and fault diagnosis), monitoring time-varying data to identify anomalies is crucial for detecting unusual behavior, potential issues, or security threats (Li and Jung, 2021; Liu et al., 2022; Chen et al., 2021; Pereira and Silveira, 2018; Li et al., 2023a). For example, in the finance field, detecting anomalous behavior in credit card data allows the prevention of theft or fraudulent transactions committed by an unauthorized party (Singh and Upadhyaya, 2012). In industrial processes, identifying anomalies helps secure safe operations, averting safety concerns and mitigating economic losses (Li et al., 2023b).
@@
In this paper, we propose a time series anomaly detection method based on INR reconstruction (TSINR for short). Specifically, we introduce a transformer-based architecture to predict the INR parameters of the given time series data. To better learn and reconstruct time series data, we design a novel form of continuous function to decompose time series (Cleveland et al., 1990; Fons et al., 2022). The designed function mainly comprises three components and individually learns the trend, seasonal, and residual information of time series. In addition, to further enhance the capability of INR to capture inter- and intra-channel information, we propose a group-based architecture to specifically learn the complex residual information. Simultaneously, we leverage a pre-trained large language model (LLM) to encode the original data to the feature domain, amplifying the fluctuations of anomalies from both time and channel dimensions and enabling INR to subsequently better differentiate normal and abnormal points. The major contributions of this paper are summarized as follows:
@@ 2.1. Time Series Anomaly Detection
Time series anomaly detection methods primarily include statistical, classic machine learning, and deep learning methods. Statistical methods rely on analyzing the statistical properties of the data to identify patterns that deviate from the expected behavior. They are valuable for their simplicity and interpretability, but have limitations in capturing complex patterns (Cleveland et al., 1990; Coluccia et al., 2013).
@@
Recently, some methods have attempted to build a universal framework for various time series data tasks, such as FPT (Zhou et al., 2024) and TimesNet (Wu et al., 2022). Among these works, models designed for prediction tasks also perform well in anomaly detection tasks, such as DLinear (Zeng et al., 2023), ETSformer (Woo et al., 2022) and LightTS (Zhang et al., 2022). These methods demonstrate outstanding performance in learning time series features, thereby it makes sense to use these methods as baselines.
@@ 2.2. Implicit Neural Representations
Current methods for learning INR parameters are primarily based on either meta-learning (Liu et al., 2023a) or feed-forward networks (Chen and Wang, 2022; Zhang et al., 2024). The key difference is that meta-learning requires a few training steps for each unseen test data, while feed-forward networks directly make predictions in a single forward pass. In this paper, we use a transformer-based architecture to generate INR parameters and it requires only a single forward in the inference phase (Chen and Wang, 2022).
@@ 3.2. Overall Workflow
Figure 2 shows the overall workflow of the proposed TSINR method. We employ a feed-forward transformer-based architecture to directly predict the whole weights of the INR of the given time series data (Chen and Wang, 2022). Unlike meta-learning based on gradient descent (Liu et al., 2023a), our method requires only a single forward step in the inference phase. Following a strategy similar to other transformer-based methods (Dosovitskiy et al., 2020; Zeng et al., 2023), the input time series data is normalized and segmented into patches. A frozen pre-trained LLM encoder is applied to map the input data into the feature domain to amplify the fluctuations of anomalies. Then the obtained features are converted to data tokens using a fully connected (FC) layer. Simultaneously, we initialize the corresponding INR tokens, which are learnable vector parameters. These data tokens and initialized INR tokens are fed together into a transformer encoder, which mainly consists of self-attention modules and feed-forward modules. In this transformer encoder, the knowledge interacts with data tokens and INR tokens. The learned INR tokens are mapped to the INR weights through FCs, denoted as FC. These INR weights form our INR continuous function, which is specifically designed for time series data. The designed function takes a batch of timestamps $\{t\}^{T}_{i=1}$ as input, implicitly learns the trends, seasonality, and residual information of the given time series data, and finally reconstructs the input signal. The details of the designed form of INR continuous function and the applied frozen pre-trained LLM encoder module are demonstrated in Section 3.3 and Section 3.4. And the anomaly criterion is introduced in Section 3.5. Relevent pseudo-codes are provided in Appendix A.
@@ 3.5. Anomaly Criterion
4. Experiments
4.1. Datasets
We use eight anomaly detection benchmarks from real-world scenarios to validate the performance of our proposed method, including seven multivariate datasets (SMD (Su et al., 2019), PSM (Abdulaal et al., 2021), SWaT (Mathur and Tippenhauer, 2016), MSL (Hundman et al., 2018), SMAP (Hundman et al., 2018), PTB-XL (Jiang et al., 2023; Wagner et al., 2020), and SKAB (Katser and Kozitsin, 2020)) and one univariate dataset (UCR (Wu and Keogh, 2021)). More details are in Appendix B.
4.2. Baselines and Experimental Settings
We compare our proposed method with 11 state-of-the-art deep learning approaches, including both general frameworks designed for time series modeling and algorithms specifically tailored for time series anomaly detection: FPT (Zhou et al., 2024), TimesNet (Wu et al., 2022), ETSformer (Woo et al., 2022), FEDformer (Zhou et al., 2022), LightTS (Zhang et al., 2022), DLinear (Zeng et al., 2023), Autoformer (Wu et al., 2021), Pyraformer (Liu et al., 2021), AnomalyTransformer (Xu et al., 2021), Informer (Zhou et al., 2021b) and Transformer (Vaswani et al., 2017). Also, we compare our TSINR method with other classical anomaly detection methods and the comparison results are in Appendix F. Additionally, we employ the commonly used metrics of precision, recall, and F1-score for evaluation.
The implementation details and the default hyper-parameters are summarized here. For a fair comparison, we only employ the classical reconstruction error across all baseline models. Also, we adopt identical data processing methods and the corresponding parameter configurations. We employ the sliding window approach and use a fixed window size of 100 for all datasets. The proportion $\gamma$ mentioned in Section 3.5 is set to 0.5 for SMD dataset, 0.1 for UCR dataset, 10 for SKAB dataset, and 1 for others. These parameters adhere to the settings of previous work (Zhou et al., 2024; Xu et al., 2021). Ablation studies on the anomaly proportion $\gamma$ are in Appendix E. For the main results, our TSINR model involves 3 global layers and 2 group layers in the residual block. And the hidden dimensions are 64 and 32 respectively. The transformer encoder has 6 blocks. We use GPT2 (Radford et al., 2019) as the pre-trained LLM encoder and 6 blocks are utilized following the same settings as in FPT (Zhou et al., 2024). The experiments are conducted using the ADAM optimizer (Kingma and Ba, 2014) with an initial learning rate of $10^{-4}$. A single NVIDIA Tesla-V100 32GB GPU is applied for each dataset. And the efficiency analysis is in Appendix H.
@@ 4.3. Main Resutls
In multivariate scenarios, we observe that despite both MSL and SMAP being collected from NASA Space Sensors, TSINR achieved significantly greater improvements on the SMAP dataset compared to other methods. This could be attributed to the presence of more point anomalies in the SMAP dataset. Point anomalies exhibit poorer continuity compared to other anomaly patterns due to their isolated nature, representing single data points that significantly deviate from the surrounding pattern. This aligns with the property of spectral bias, making our model more sensitive to point anomalies, thereby achieving greater improvements on the SMAP dataset. Meanwhile, our approach shows moderate performance on the SWAT dataset. This is due to the poorer continuity of the SWaT dataset, which affects the fitting ability of TSINR. More detailed analysis can be found in Appendix C.
@@
The situation is more complex in the UCR dataset. The UCR dataset comprises 250 univariate sub-datasets, and we report the average scores. The results demonstrate that our approach still outperforms other methods in overall performance by a significant margin. These sub-datasets originate from various domains, demonstrating the generalization capability of our method. In addition, it proves the strong ability of TSINR for modeling and identifying anomalies in univariate scenarios and underscores the importance of the spectral bias constraint for anomaly detection. We compile the performance of models in each domain of the UCR dataset, with the complete experimental results presented in Appendix G.
@@ 4.4.1. Analysis of the Decomposition Components and the Group-based Architecture
The main purpose of the decomposition components is to extract the unique trend and seasonal information of the time series data. The results in Table 2 indicate that capturing these distinctive features significantly enhances the capability for anomaly detection. In addition, the group-based architecture is designed to enhance the representational capacity of INR for multivariate data. Experimental results indicate an improvement in the capability for anomaly detection when employing the proposed group-based architecture. This is because modeling multiple variables and capturing both inter- and intra-channel information with a simple continuous function, which only consists of fully-connected layers, is challenging. Our approach addresses this by dividing the variables into several groups and applying independently fully-connected layers in different groups, thereby reducing the number of variables each group needs to model and improving the representational capacity. The global layers extract the inter-channel information, while the group layers selectively focus on detailed information for specific channels. This enhances the representational capability for each variable without losing any knowledge. Detailed ablation studies on the number of groups are left in Appendix D.
@@ 5. Conclusion
Appendix A Algorithm
Pseudo-codes of the main components of TSINR are presented in 1. All experiments are implemented based on PyTorch. The proposed form of INR continuous function takes the timestamp $t$ as input, which individually passes through three blocks: trend, seasonal, and residual, to predict the corresponding time series information. Then, the predictions from these blocks are summed up to obtain the reconstructed time series data as shown in Eq. 1. For the trend and seasonal components, we use two single-layer MLPs to fit the trend information in Eq. 2 and the periodic information in Eq. 3 separately. For the residual component, we propose a group-based architecture to learn intricate residual information. A 5-layer MLP is utilized for learning complex residual information in Eq. 4 and Eq. 5, comprising 3 global layers and 2 group layers. Finally, we use a transformer-based architecture to generate the INR parameters. The original data is encoded by a pre-trained LLM encoder and the obtained features are passed into the transformer encoder as data tokens. The INR tokens are learned through transformer encoders and further transformed into parameters for the MLPs in the trend, seasonal, and residual blocks through the post functions.
Algorithm 1 TSINR Anomaly Detection Algorithm
1:Input: A batch of time serial data $X\in\mathbb{R}^{B\times d\times T}$, trained initialized INR token $I_{init}\in\mathbb{R}^{B\times N_{I}\times C}$, the timestamp $C\in R^{B\times T}$, the anomaly threshold $\delta$
2:Output: The predicted anomaly label $Y\in[0,1]^{B\times T}$
3:Normalize and patch $X$ to data feature $D\in\mathbb{R}^{B\times N_{d}\times C}$
4:Concatenate $D$ and $I_{init}$ to feature $F\in\mathbb{R}^{B\times(N_{d}+N_{I})\times C}$
5:Encode $F$ with LLM encoder and learnable Transformer encoder to get processed feature $F^{\prime}\in\mathbb{R}^{B\times(N_{d}+N_{I})\times C}$
6:Obtain the data-dependant INR token $I^{\prime}\in\mathbb{R}^{B\times N_{I}\times C}$ by choosing the corresponding part of $F^{\prime}$
7:Using FC layer to convert I’ to the parameter $\theta$ of MLP and form the INR continuous function $f_{\theta}$
8:Query the corresponding value of the timestamp $C$ by $X^{\prime}=f_{\theta}(C)$ to get the reconstruct the time serial data $X^{\prime}\in\mathbb{R}^{B\times d\times T}$
9:Calculate the Anomaly Score $S\in R^{B\times T}$by measuring the point-wise difference between the input data $X$ and the reconstructed data $X^{\prime}$ as in equation (7).
10:Determine whether the point is abnormal or normal with the threshold $\delta$ as in equation (8) and get the predicted anomaly label $Y\in[0,1]^{B\times T}$
Appendix B Dataset Description
We conduct comprehensive experiments on seven multivariate datasets and one univariate dataset to validate the effectiveness of our proposed TSINR model.
Multivariate Datasets: (1) SMD (Server Machine Dataset) is a novel 5-week-long dataset collected from a large Internet company, comprising data from 28 distinct machines with 38 dimensions (Su et al., 2019). (2) PSM (Pooled Server Metrics) is sourced internally from various application server nodes at eBay with 25 dimensions and consists of 13 weeks of training data and 8 weeks of testing data (Abdulaal et al., 2021). (3) SWaT (Secure Water Treatment) is collected from an actual industrial water treatment plant and involves 51 dimensions (Mathur and Tippenhauer, 2016). (4) MSL (Mars Science Laboratory) and (5) SMAP (Soil Moisture Active Passive) are collected by spacecraft monitoring systems from NASA and comprises 55 and 25 dimensions respectively (Hundman et al., 2018, 2018). (6) PTB-XL includes clinical 12-lead Electrocardiography (ECG) that is 10 seconds in length for each patient (Jiang et al., 2023; Wagner et al., 2020). (7) SKAB is collected from the sensors installed on the testbed and includes 8 dimensions (Katser and Kozitsin, 2020).
Univariate Dataset: (8) UCR is a new benchmark introduced by multi-dataset time series anomaly detection competition in 2021 (Wu and Keogh, 2021). It encompasses 250 univariate sub-datasets, spanning diverse domains. Synthetic but highly plausible anomalies are created and each time series data in UCR has only one anomaly.
Anomaly ratio refers to the true abnormal proportion of the whole dataset while anomaly proportion indicates the abnormal proportion the user decided to choose the corresponding threshold. Anomaly proportion and window size are hyper-parameters. We follow the experimental settings of previous studies (Zhou et al., 2024; Xu et al., 2021).
Appendix C Discussion on Continuity of Multivariable Datasets
Figure 8 is a kernel density estimation plot, which is a method for visualizing the distribution of observations in a dataset. It can be observed that the SWAT dataset has the most peaks, and there is a lack of transition points between the peaks, indicating that the data changes significantly, are more high-frequency, and have poor continuity. This limitation adversely impacts the fitting capability of TSINR, consequently affecting the performance of anomaly detection.

### v3

@@ Abstract.
copyright: acmlicenseddoi: 3690624.3709266conference: the 31th ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadaccs: Computing methodologies Neural networksccs: Mathematics of computing Time series analysis
@@ 1. Introduction
Time series anomaly detection, which aims to identify unusual patterns or events across a sequence of data points collected over time (Li and Jung, 2023), has attracted a lot of attention recently. In many fields (e.g., finance, healthcare, manufacturing, and fault diagnosis), monitoring time-varying data to identify anomalies is crucial for detecting unusual behavior, potential issues, or security threats (Wang et al., 2018; Jie et al., 2024; Wang et al., 2019; Wu et al., 2024; Li et al., 2023a). For example, in the finance field, detecting anomalous behavior in credit card data allows the prevention of theft or fraudulent transactions committed by an unauthorized party (Singh and Upadhyaya, 2012). In industrial processes, identifying anomalies helps secure safe operations, averting safety concerns and mitigating economic losses (Li et al., 2023b).
@@
In this paper, we propose a time series anomaly detection method based on INR reconstruction (TSINR for short). Specifically, we introduce a transformer-based architecture to predict the INR parameters of the given time series data. To better learn and reconstruct time series data, we design a novel form of continuous function to decompose time series (Cleveland et al., 1990; Fons et al., 2022). The designed function mainly comprises three components and individually learns the trend, seasonal, and residual information of time series. In addition, to further enhance the capability of INR to capture inter- and intra-channel information, we propose a group-based architecture to specifically learn the complex residual information. Simultaneously, we leverage a pre-trained large language model (LLM) to encode the original data to the feature domain. This encoding process particularly amplifies the fluctuations of anomalies across both the time and channel dimensions. By doing so, we enable TSINR to more effectively distinguish between normal and abnormal data points. The major contributions of this paper are summarized as follows:
@@ 2.1. Time Series Anomaly Detection
Time series anomaly detection methods primarily include statistical, classic machine learning, and deep learning methods. Statistical methods rely on analyzing the statistical properties of the data to identify patterns that deviate from the expected behavior. They are valuable for their simplicity and interpretability, but have limitations in capturing complex patterns (Cleveland et al., 1990; Jie et al., 2024).
@@
Recently, several methods have been proposed to establish a universal framework that can effectively address a wide range of time series data tasks, such as FPT (Zhou et al., 2024) and TimesNet (Wu et al., 2022). Among these works, models designed for prediction tasks also perform well in anomaly detection tasks, such as DLinear (Zeng et al., 2023), ETSformer (Woo et al., 2022) and LightTS (Zhang et al., 2022). These methods demonstrate outstanding performance in learning time series features, thereby it makes sense to use these methods as baselines.
@@ 2.2. Implicit Neural Representations
Current methods for learning INR parameters predominantly rely on two main approaches: meta-learning (Liu et al., 2023a) and feed-forward networks (Chen and Wang, 2022; Zhang et al., 2024). The primary distinction between these approaches lies in how they handle test data. Meta-learning techniques are designed to quickly adapt to unseen test data by requiring only a few training steps or fine-tuning iterations for each new input. This allows for efficient generalization to novel tasks with minimal data. In contrast, feed-forward networks directly generate predictions in a single forward pass, leveraging pre-trained parameters without the need for task-specific adaptation during inference. In this paper, we use a transformer-based architecture to generate INR parameters and it requires only a single forward in the inference phase (Chen and Wang, 2022).
@@ 3.2. Overall Workflow
Figure 2 shows the overall workflow of the proposed TSINR method. We employ a feed-forward transformer-based architecture to directly predict the whole weights of the INR of the given time series data (Chen and Wang, 2022). Unlike meta-learning based on gradient descent (Liu et al., 2023a), our method requires only a single forward step in the inference phase. Following a strategy similar to other transformer-based methods (Dosovitskiy et al., 2020; Zeng et al., 2023), the input time series data is normalized and segmented into patches. A frozen pre-trained LLM encoder is applied to map the input data into the feature domain to amplify the fluctuations of anomalies. Then the obtained features are converted to data tokens using a fully connected (FC) layer. Simultaneously, we initialize the corresponding INR tokens, which are learnable vector parameters. These data tokens and initialized INR tokens are fed together into a transformer encoder, which mainly consists of self-attention modules and feed-forward modules. In this transformer encoder, the knowledge interacts with data tokens and INR tokens. The learned INR tokens are mapped to the INR weights through FCs, denoted as FC. These INR weights form our INR continuous function, which is specifically designed for time series data. The designed function takes a batch of timestamps $\{t\}^{T}_{i=1}$ as input, implicitly learns the trends, seasonality, and residual information of the given time series data, and finally reconstructs the input signal. The details of the designed form of INR continuous function and the applied frozen pre-trained LLM encoder module are demonstrated in Section 3.3 and Section 3.4. The anomaly criterion is introduced in Section 3.5.
@@ 3.5. Anomaly Criterion
4. Experiments
4.1. Datasets
We use eight anomaly detection benchmarks from real-world scenarios to validate the performance of our proposed method, including seven multivariate datasets (SMD (Su et al., 2019), PSM (Abdulaal et al., 2021), SWaT (Mathur and Tippenhauer, 2016), MSL (Hundman et al., 2018), SMAP (Hundman et al., 2018), PTB-XL (Jiang et al., 2023; Wagner et al., 2020), and SKAB (Katser and Kozitsin, 2020)) and one univariate dataset (UCR (Wu and Keogh, 2021)).
4.2. Baselines and Experimental Settings
We compare our proposed method with 11 state-of-the-art deep learning approaches, including both general frameworks designed for time series modeling and algorithms specifically tailored for time series anomaly detection: FPT (Zhou et al., 2024), TimesNet (Wu et al., 2022), ETSformer (Woo et al., 2022), FEDformer (Zhou et al., 2022), LightTS (Zhang et al., 2022), DLinear (Zeng et al., 2023), Autoformer (Wu et al., 2021), Pyraformer (Liu et al., 2021), AnomalyTransformer (Xu et al., 2021), Informer (Zhou et al., 2021b) and Transformer (Vaswani et al., 2017). The commonly used metrics of precision (P), recall (R), and F1-score are employed for evaluation. Additionally, we report threshold-free measurements, including the Area Under the Curve (AUC) and Volume Under the Surface (VUS) (Paparrizos et al., 2022), which are provided in the Appendix A.
The implementation details and the default hyper-parameters are summarized here. For a fair comparison, we only employ the classical reconstruction error across all baseline models. Also, we adopt identical data processing methods and the corresponding parameter configurations. We employ the sliding window approach and use a fixed window size of 100 for all datasets. The proportion $\gamma$ mentioned in Section 3.5 is set to 0.5 for SMD dataset, 0.1 for UCR dataset, 10 for SKAB dataset, and 1 for others. These parameters adhere to the settings of previous work (Zhou et al., 2024; Xu et al., 2021). Ablation studies on the anomaly proportion $\gamma$ are in Appendix B. For the main results, our TSINR model involves 3 global layers and 2 group layers in the residual block. And the hidden dimensions are 64 and 32 respectively. The transformer encoder has 6 blocks. We use GPT2 (Radford et al., 2019) as the pre-trained LLM encoder and 6 blocks are utilized following the same settings as in FPT (Zhou et al., 2024). The experiments are conducted using the ADAM optimizer (Kingma and Ba, 2014) with an initial learning rate of $10^{-4}$. A single NVIDIA Tesla-V100 32GB GPU is applied for each dataset. And the efficiency analysis is in Appendix C.
@@ 4.3. Main Resutls
In multivariate scenarios, we observe that despite both MSL and SMAP being collected from NASA Space Sensors, TSINR achieved significantly greater improvements on the SMAP dataset compared to other methods. This could be attributed to the presence of more point anomalies in the SMAP dataset. Point anomalies exhibit poorer continuity compared to other anomaly patterns due to their isolated nature, representing single data points that significantly deviate from the surrounding pattern. This aligns with the property of spectral bias, making our model more sensitive to point anomalies, thereby achieving greater improvements on the SMAP dataset. Meanwhile, our approach shows moderate performance on the SWAT dataset. This is due to the poorer continuity of the SWaT dataset, which affects the fitting ability of TSINR.
The situation is more complex in the UCR dataset. The UCR dataset comprises 250 univariate sub-datasets from various domains and we report the average scores to provide a comprehensive assessment. The results demonstrate that our approach still outperforms other methods in overall performance by a significant margin. These sub-datasets originate from various domains, demonstrating the generalization capability of our method. In addition, it proves the strong ability of TSINR for modeling and identifying anomalies in univariate scenarios and underscores the importance of the spectral bias constraint for anomaly detection.
@@
Table 3. Analysis of the decomposition components. The MSE and F1 scores (*/*) are reported and the best results are in Bold.
Decomposition Synthetic Trend Synthetic Seasonal SMD PSM MSL SKAB
0.56/71.15 0.44/25.97 1.21/80.10 0.26/93.08 2.23/82.28 1.23/94.01
0.09/100.00 0.01/100.00 0.99/81.76 0.21/94.04 1.81/84.74 0.84/94.11
@@ 4.4.1. Analysis of the Decomposition Components and the Group-based Architecture
The main purpose of the decomposition components is to extract the unique trend and seasonal information of the time series data. The results in Table 2 indicate that capturing these distinctive features significantly enhances the capability for anomaly detection. To further explain the effectiveness of the trend and seasonal components in INR continuous function, we conduct ablation experiments using both synthetic and real-world datasets. The synthetic trend and seasonal datasets are generated by (Lai et al., 2021), while the real-world datasets include SMD, PSM, MSL, and SKAB. As shown in Table 3 and Figure 4, incorporating trend and seasonal components significantly improves both reconstruction performance and anomaly detection capability. In contrast, omitting these components results in inadequate data fitting in certain cases, thereby hindering the ability to detect anomalies. Also, we showcase the detection of a non-spike anomaly segment in Figure 5, highlighting the role of trend and seasonal components in identifying such anomalies. The anomaly, characterized by subtle deviations from the expected behavior, is effectively captured, demonstrating the ability of TSINR to detect anomalies that do not exhibit abrupt or spike-like changes. This emphasizes the robustness of the model in handling different types of anomalies in real-world datasets like PSM.
Figure 4. The visual analysis of the decomposition components on synthetic trend and seasonal datasets.
Figure 5. The visual analysis of the non-spike anomaly segment in the real-world PSM dataset.
In addition, the group-based architecture is designed to enhance the representational capacity of INR for multivariate data. Experimental results indicate an improvement in the capability for anomaly detection when employing the proposed group-based architecture. This is because modeling multiple variables and capturing both inter- and intra-channel information with a simple continuous function, which only consists of fully-connected layers, is challenging. Our approach addresses this by dividing the variables into several groups and applying independently fully-connected layers in different groups, thereby reducing the number of variables each group needs to model and improving the representational capacity. The global layers extract the inter-channel information, while the group layers selectively focus on detailed information for specific channels. This enhances the representational capability for each variable without losing any knowledge. Detailed ablation studies on the number of groups are left in Appendix D.
@@ 5. Conclusion
6. Acknowledgments
This work was supported by the National Natural Science Foundation of China (Grant No. 62202422), the National Key R&D Program of China (Grant No. 2022ZD0160703), the National Natural Science Foundation of China (Grant Nos. 62276230 and 62372408), Zhejiang Provincial Natural Science Foundation of China (Grant No. LDT23F02023F02), and Shanghai Artificial Intelligence Laboratory.
@@ 6. Acknowledgments
Appendix A Additional Results
In addition to the commonly used F1-score, which is threshold-dependent, we evaluate TSINR using threshold-free AUC and VUS scores. These metrics provide a more comprehensive assessment of model performance, as they do not rely on a specific decision threshold. The results in Table demonstrate that TSINR consistently outperforms competing methods across all datasets, highlighting its robustness and superior capability in anomaly detection tasks.
Appendix B Study on the Anomaly Proportion
Appendix C Efficiency Analysis